OpenStudy (anonymous):

You are getting ready to move and have asked some friends to help. For lunch, you guy the following sandwiches at the local deli for $30; six tuna sandwiches and six turkey sandwiches. Later at night, you again buy 4 tuna sandwiches, and 8 turkey ones for $30. 60. What is the price of each sandwich?

OpenStudy (anonymous):

let cost of one tuna sw = x let cost of one turkey sw = y then u get two equations 6x + 6y = 30 4x + 8y = 30.60 solve them to get your answer

OpenStudy (anonymous):

thank you!! :)

OpenStudy (anonymous):

if you have time, can u help me with the last problem i don't get? :)

OpenStudy (anonymous):

u r welcome....☺ you know how to solve it, right???

OpenStudy (anonymous):

go ahead..

OpenStudy (anonymous):

yes i do know how to solve it:) & You have a cable plan that costs $39 a month for a basic plan plus one movie channel. Your friend has the same basic plan plus 2 movie channels for 45.50. What is the basic plan charge that you both pay?

OpenStudy (anonymous):

let the basic plan cost be = x let the cost of movie channel be = y your cost : x + y = 39 friend's cost : x + 2y = 45.50